# version 0.75

### v4

@@ 7.3.1 Convexity of the losses
To conclude the analysis, We first show that the loss functions are convex w.r.t. to the variables ${M}_{1:h}$. This follows since the states and the controls are linear transformations of the variables.
@@
Since the cost $c_{t}$ is a convex function w.r.t. its arguments, we simply need to show that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{1:h}$. The state is given by
@@ 7.3.2 Loss functions with memory
The actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{x}_{t}(M_{1:h})$, but rather on the true state $\mathbf{x}_{t}$, which in turn depends on different parameters $M_{1:h}^{i}$ for various historical times $i<t$. Nevertheless, $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{1:h})$, as shown next.
@@
$=\frac{Ch\kappa D}{\sqrt{T}}\sum_{i=1}^{\infty}i(1-\delta)^{i-1}$ $\sum_{i=1}^{\infty}i(1-\delta)^{i}=\frac{1}{\delta^{2}}$
$\leq\frac{C\gamma^{2}hD}{\sqrt{T}}.$
@@ 7.4 Bibliographic Remarks
The bulk of the material of this chapter is from [ABH+19]. The work of [ABH+19] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions. Follow up work by [AHS19] achieves logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise. Logarithmic regret algorithms were derived for online nonstochastic control in [Sim20, FS20].
Online nonstochastic control with known dynamics was further extended in several directions. Control with bandit feedback for the costs was considered in [GHH20]. More refined regret metrics applicable to changing dynamics were studied in [GHM20, MGSH21, ZWZ22, Sin22]. More extensions suitable for changing environments were studied in [BW22].
[LDL21] studies algorithms that balance regret minimization while adhering to constraints. An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHMS21]. Deluca [GHM20], a control library, provides benchmarked implementations of many such algorithms.
Online nonstochastic control is an extension of the framework of online convex optimization [Haz16] when applied to policies, and thus to loss functions with memory [AHM13]. A similar regret notion was implicit in the work of [EDKM04]. The regret metric when applied to policies is different from the counterfactual notion of policy regret as defined by [ADT12].
An alternate performance metric for online control, which is also nonstochastic, is competitive ratio. This objective tracks the cost of the online controller and compares it to the best policy in hindsight. Control methods that minimize competitive ratio were initiated in [GW19]. This approach was extended in a series of papers [GLSW19, SLC+20, GH21]. Thus far the success of this approach was limited to fully observed linear dynamical systems, and further to quadratic costs when achieving the optimal competitive ratio. [GASH22] show that regret minimizers in the non-stochastic setting automatically achieve a near-optimal competitive ratio.
A precursor to online nonstochastic control is a line of work, beginning with [AYBK14, CHK+18], focused on the control of stochastic systems under sequentially revealed adversarial costs functions. See [CK20, CCK22] for further developments.
@@ 7.5 Exercises
7.5. In this exercise we prove and extend Theorem 7.2 for linear time-varying systems that are stabilizable, rather than stable.
@@
In the previous chapters, we focused on control algorithms that assume knowledge of the system dynamics. However, in many real-world applications, system parameters are either unknown or change over time. This necessitates an approach where the controller must simultaneously learn the system model while making control decisions. This problem, known as system identification, is fundamental to adaptive control, reinforcement learning, and modern data-driven methods for control.
@@ 8.1 The Role of System Identification in Online Control
  • Balancing exploration and performance: The controller must explore sufficiently to estimate system parameters accurately while ensuring stable performance, a tradeoff that directly impacts regret minimization.
@@ 8.2 A Nonstochastic Approach to System Identification
  • Computational tractability in identification: Classical system identification methods often involve solving nonconvex optimization problems, making them computationally demanding. Nonstochastic approaches frequently leverage convex relaxation techniques to sidestep intractability.
The problem of online system identification is closely related to online convex optimization and sequential decision-making. Many regret-minimization techniques developed for online learning can be adapted to system identification settings. A key distinction, however, is that in control settings, the learned system model directly affects future decisions, creating a coupling between estimation and control that does not typically exist in standard online learning formulations.
@@ 8.3 Nonstochastic System Identification
The method below for achieving this learns the system matrices by using a random control. Specifically, at each step the control $\mathbf{u}_{t}$ will be randomly selected from a Rademacher distribution, as per Algorithm 5.
@@
  Define $\widehat{C_{0}}=(\widehat{B},\dots\widehat{A^{k-1}B})$, $\widehat{C_{1}}=(\widehat{A^{1}B},\dots\widehat{A^{k}B})$, and return $\widehat{A},\widehat{B}$ as
@@
and the crucial fact that the controls are i.i.d. Thus, we can create an unbiased estimator for this matrix, and using many samples get an increasingly accurate estimate.
@@
By this display, we already have $\|B-\widehat{B}\|\leq\varepsilon$. We will make use of this observation to establish the proximity of $\widehat{A}$ and $A$, whose proof is left as Exercise 8.7.
Note that $\widehat{A}$ does not necessarily satisfy $\widehat{A}\widehat{C_{0}}=\widehat{C_{1}}$. Hence we require the following lemma.
@@
Let $X,X+\Delta X$ be the solutions to the linear system of equations $XA=B$ and $(X+\Delta X)(A+\Delta A)=B+\Delta B$. Then, as long as $\|\Delta A\|\leq\sigma_{\min}(A)$, we have
@@
We now use this lemma on the normal equations, i.e. $AC_{0}C_{0}^{\top}=C_{1}C_{0}^{\top}$ and $\widehat{A}\widehat{C_{0}}\widehat{C_{0}}^{\top}=\widehat{C_{1}}\widehat{C_{0}}^{\top}$. First, observe that for any $\varepsilon<1$
@@
$\|A-\widehat{A}\|\leq\frac{12\gamma^{2}k}{\sigma_{\min}(A)}\varepsilon,$
as long as $3k\kappa\varepsilon/\delta<\sigma_{\min}(C_{0})$. Finally, note that $(k,\kappa)$-strong controllability implies $\sigma_{\min}(C_{0})\geq 1/\kappa$. ∎
@@ 8.5 Summary and Key Takeaways
  • The Exploration-Exploitation Tradeoff A fundamental challenge in online system identification is balancing exploration, which enables learning the system dynamics, and exploitation, which minimizes control cost. Too much exploration leads to high immediate costs, while too much exploitation can result in poor long-term learning and suboptimal control performance. The algorithms in this chapter address this by incorporating controlled perturbations in the control inputs, ensuring sufficient exploration while keeping regret low. However, the tradeoff results in a regret of $O(T^{2/3})$, rather than $O(\sqrt{T})$ we had in the last chapter.
@@
  • Regret Minimization as a Learning Metric Unlike traditional identification methods that aim for asymptotic convergence, online nonstochastic identification is evaluated in terms of regret, measuring performance relative to the best fixed model in hindsight.
@@ 8.5.1 Looking Ahead
This part of the book considered the online nonstochastic control problem, the regret metric, and online learning techniques applied to control. We have thus far considered the case where the state is observable to the controller, an assumption which doesn’t always hold.
In the next parts of the book, we consider the setting of partial observability, where the system state is unknown, but an observed signal is available. We start with the problem of learning, and move onto online control.
@@ 8.6 Bibliographic Remarks
Controlling linear dynamical systems given unknown or uncertain system parameters has long been studied in adaptive control; see the text of [SBB90] for example. However, the results from adaptive control typically concern themselves with stability, and even when discussing implications for optimality present asymptotic bounds.
In contrast, much of the work at the intersection of machine learning and control is focussed on precise non-asymptotic bounds on sample complexity or regret. [Fie97] which studies the discounted LQR problem under stochastic noise is the earliest example that furnishes a precise sample complexity bound. Similarly, the early work of [AYS11] uses an optimism-based approach to achieve an optimal regret in terms of the dependence on horizon. This regret bound was later improved for sparse systems in [IJR12]. [DMM+18, DMM+20] established regret and sample complexity results for the general stochastic LQR formulation with a polynomial dependence on natural system parameters. [FGKM18] studies the efficacy of policy-gradient-based approaches for synthesizing controllers for the LQR problem. The work of [MTR19] establishes that simple explore-then-commit strategies achieve the optimal dependence of regret on horizon, with [SF20] proving that such approaches are optimal in terms of their dependence on dimension of the system too. [CKM19] also gave a first optimal-in-horizon polynomial regret result simultaneously with the former work albeit using an optimism-based approach. The result of [PH20] gives a rate-optimal regret bound for unknown systems with convex costs.
In terms of nonstochastic control, the method-of-moments system identification procedure and the corresponding analysis presented here appeared in [HKS20], which also gave the first end-to-end regret result for nonstochastic control starting with unknown system matrices. Even earlier, [SBR19] presented a regression procedure for identifying the Markov operators under nonstochastic noise. Both of these approaches have roots in the classical Ho-Kalman [HK66] system identification procedure, for which the first nonasymptotic sample complexity, although for stochastic noise, were furnished in [OO19]. For strongly convex costs, logarithmic regret was established in [Sim20]. Nonstochastic control of unknown, unstable systems without any access to prior information, also called blackbox control, was addressed in [CH21].
@@ 8.7 Exercises
In previous chapters, we focused on control in dynamical systems, often assuming that the system state was fully observable and that system dynamics were either known or could be estimated through system identification. In many real-world scenarios, however, these assumptions do not hold. Instead, an agent may only have access to partial observations of the system and must learn to make accurate predictions without directly influencing the system’s evolution.
This chapter focuses on learning in a dynamical system, which means predicting future system outputs based on past observations while the underlying system parameters remain unknown. Unlike control, where an agent actively influences system trajectories to minimize a given cost, learning in this setting is a passive task—the goal is to accurately forecast future outputs without modifying the system’s behavior.
The ability to predict future states and outputs of an unknown system is fundamental to many domains, including time-series forecasting, financial modeling, sensor networks, autonomous systems, and most recently large language models. Moreover, learning a predictive model is often a key stepping stone toward solving more complex problems, such as control under partial observability or reinforcement learning in dynamical systems. In this chapter, we study online learning algorithms for prediction in unknown linear dynamical systems, with an emphasis on regret minimization.
@@ 9.1 Learning in Dynamical Systems
The term learning in dynamical systems can refer to a variety of tasks, including state estimation, system identification, and policy learning. In this chapter, we focus on the specific task of prediction, where the objective is to estimate future outputs of a dynamical system based on past observations. Unlike classical system identification, which seeks to recover the system parameters explicitly, the learning approach taken here aims to optimize predictions directly.
@@
For the remainder of this chapter, we consider a partially observable linear time invariant dynamical system described by the following equations. The extention of the material in this chapter to time varying linear dynamical systems is left as exercise 9.7.
@@
Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ is the observed control input, and $\mathbf{y}_{t}$ is the measured system output. The terms $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ account for process and observation noise, respectively. A crucial challenge in this setting is that the system matrices $A,B,C$ are unknown to the learner.
The learner’s objective is to construct an efficient prediction rule that estimates $\mathbf{y}_{t}$ based on past observations, despite the unknown system parameters.
@@ 9.2 Online Learning of Dynamical Systems
In an online learning setting, the learner sequentially observes data and makes predictions in real time. Formally, the process unfolds as follows:
@@
    The learner produces a prediction $\hat{\mathbf{y}}_{t}$ for the next output.
@@
This regret-based framework allows us to compare different classes of predictors and design learning algorithms that adapt efficiently to unknown dynamical systems.
@@ 9.3 Classes of Prediction Rules
We now explore what constitutes a reasonable prediction rule for the output of a linear dynamical system. However, before considering the prediction rules themselves, we discuss how to compare classes of such rules. This is analogous to the definitions and discussion in chapter 6 where we discuss comparison between policy classes for control. However, the definitions and conclusions for prediction are simpler.
@@
Proof of this lemma is left as an exercise to the reader at the end of this chapter. We now proceed to consider specific prediction rules.

### v5

@@ 7.3.1 Convexity of the Losses
To conclude the analysis, we first show that the loss functions are convex with respect to the variables ${M}_{1:h}$. This follows since the states and the controls are linear transformations of the variables.
@@
Since the cost $c_{t}$ is a convex function with respect to its arguments, we simply need to show that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{1:h}$. The state is given by
@@ 7.3.2 Loss Functions with Memory
The actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{x}_{t}(M_{1:h})$, but rather on the true state $\mathbf{x}_{t}$, which in turn depends on different parameters $M_{1:h}^{i}$ for various historical times $i<t$. However, $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{1:h})$, as shown next.
@@
$=\frac{Ch\kappa D}{\sqrt{T}}\sum_{i=1}^{\infty}i(1-\delta)^{i-1}$
$\leq\frac{C\gamma^{2}hD}{\sqrt{T}},$
where on the second last line we use $\sum_{i=1}^{\infty}i(1-\delta)^{i}=\frac{1}{\delta^{2}}$.
@@ 7.4 Bibliographic Remarks
Online nonstochastic control is an extension of the framework of online convex optimization [Haz16], as it is applied to policies, and thus to loss functions with memory [AHM13].
The bulk of the material in this chapter is from [ABH+19]. The work of [ABH+19] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions.
The follow-up work by [AHS19] achieves a logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise. Logarithmic regret algorithms for the full setting of online nonstochastic control with strongly convex costs were obtained in [Sim20, FS20, CCK20].
Online nonstochastic control with known dynamics was further extended in several directions. The task of control with bandit feedback for the costs was considered in [GHH20]. More refined regret metrics applicable to changing dynamics were studied in [GHM20, MGSH21, ZWZ22, Sin22]. More extensions suitable for changing environments were studied in [BW22]. Online nonstochastic control of population dynamics was studied in [GHL+24].
[LDL21] studies algorithms that balance regret minimization while adhering to constraints. An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHMS21]. Various open-source software packages, such as [GHM20], provide benchmarked implementations of many such algorithms.
An alternate performance metric for online control, which is also nonstochastic, is the competitive ratio. This objective tracks the cost of the online controller and compares it to the best policy in hindsight. Control methods that minimize the competitive ratio were initiated in [GW19]. This approach was extended in a series of papers [GLSW19, SLC+20, GH21]. Thus far, the success of this approach was limited to fully observed linear dynamical systems and further to quadratic costs when achieving the optimal competitive ratio. [GASH22] show that regret minimizers in the non-stochastic setting automatically achieve a near-optimal competitive ratio.
A precursor to online nonstochastic control is a line of work, beginning with [AYBK14], focused on the control of stochastic systems under sequentially revealed adversarial costs functions. See [CK20, CCK22] for further developments.
@@ 7.5 Exercises
7.5. In this exercise, we prove and extend Theorem 7.2 for linear time-varying systems that are stabilizable, rather than stable.
@@
In the previous chapters, we focused on control algorithms that assume knowledge of the system dynamics. However, in many real-world applications, system parameters are unknown or change over time. This necessitates an approach in which the controller must simultaneously learn the system model while making control decisions. This problem, known as system identification, is fundamental to adaptive control, reinforcement learning, and modern data-driven methods for control.
@@ 8.1 The Role of System Identification in Online Control
  • Balancing exploration and performance: The controller must explore sufficiently to estimate system parameters accurately while ensuring stable performance, a trade-off that directly impacts regret minimization.
@@ 8.2 A Nonstochastic Approach to System Identification
  • Computational tractability in identification: Classical system identification methods often involve solving non-convex optimization problems, which makes them computationally demanding. Nonstochastic approaches frequently leverage convex relaxation techniques to sidestep intractability.
The problem of online system identification is closely related to online convex optimization and sequential decision making. Many regret minimization techniques developed for online learning can be adapted to system identification settings. However, a key distinction is that, in control settings, the learned system model directly affects future decisions, creating a coupling between estimation and control that does not typically exist in standard online learning formulations.
@@ 8.3 Nonstochastic System Identification
The method below to achieve this learns the system matrices by using a random control. Specifically, at each step the control $\mathbf{u}_{t}$ will be randomly selected from a Rademacher distribution, according to Algorithm 5.
@@
  Define $\widehat{C_{0}}=(\widehat{B},\dots\widehat{A^{k-1}B})$, $\widehat{C_{1}}=(\widehat{AB},\dots\widehat{A^{k}B})$, and return $\widehat{A},\widehat{B}$ as
@@
and the crucial fact that the controls are i.i.d. Thus, we can create an unbiased estimator for this matrix, and using many samples, we obtain an increasingly accurate estimate.
@@
By this display, we already have $\|B-\widehat{B}\|\leq\varepsilon$. We will make use of this observation to establish the proximity of $\widehat{A}$ and $A$, using the following lemma, the proof of which is left as Exercise 8.7.
@@
Let $X,X+\Delta X$ be the solutions to the linear system of equations $XA=B$ and $(X+\Delta X)(A+\Delta A)=B+\Delta B$. Then, as long as $\|\Delta A\|<\sigma_{\min}(A)$, we have
@@
Note that $\widehat{A}$ does not necessarily satisfy $\widehat{A}\widehat{C_{0}}=\widehat{C_{1}}$. We now use the above lemma on the normal equations, that is, on $AC_{0}C_{0}^{\top}=C_{1}C_{0}^{\top}$ and $\widehat{A}\widehat{C_{0}}\widehat{C_{0}}^{\top}=\widehat{C_{1}}\widehat{C_{0}}^{\top}$. First, observe that for any $\varepsilon<1$
@@
$\|A-\widehat{A}\|\leq\frac{12\gamma^{2}k}{\sigma_{\min}(C_{0}C_{0}^{\top})}\varepsilon,$
as long as $6k\gamma\varepsilon<\sigma_{\min}(C_{0}C_{0}^{\top})$. Finally, note that $(k,\kappa)$-strong controllability implies $\sigma_{\min}(C_{0}C_{0}^{\top})\geq 1/\kappa$. ∎
@@ 8.5 Summary and Key Takeaways
  • The Exploration-Exploitation Tradeoff A fundamental challenge in online system identification is balancing exploration, which enables learning the system dynamics, and exploitation, which minimizes control cost. Too much exploration leads to high immediate costs, while too much exploitation can result in poor long-term learning and suboptimal control performance. The algorithms in this chapter address this by incorporating controlled perturbations in the control inputs, ensuring sufficient exploration while keeping regret low. However, the trade-off results in a regret of $O(T^{2/3})$, rather than $O(\sqrt{T})$ we had in the last chapter.
@@
  • Regret Minimization as a Learning Metric Unlike traditional identification methods that aim at asymptotic convergence, online nonstochastic identification is evaluated in terms of regret, measuring the performance relative to the best fixed model in hindsight.
@@ 8.5.1 Looking Ahead
This part of the book considered the online nonstochastic control problem, the regret metric, and online learning techniques applied to control. We have thus far considered the case where the state is observable to the controller, an assumption which does not always hold.
In the next parts of the book, we consider the setting of partial observability, where the system state is unknown, but an observed signal is available. We start with the problem of learning and move onto online control.
@@ 8.6 Bibliographic Remarks
Controlling linear dynamical systems given unknown or uncertain system parameters has long been studied in adaptive control; see the text of [SBB90] for example. However, the results from adaptive control typically concern themselves with stability and, even when discussing implications for optimality, present asymptotic bounds.
In contrast, much of the work at the intersection of machine learning and control is focussed on precise non-asymptotic bounds on sample complexity or regret. [Fie97] which studies the discounted LQR problem under stochastic noise is the earliest example that provides a precise sample complexity bound. Similarly, the early work of [AYS11] uses an optimism-based approach to achieve optimal regret in terms of dependence on horizon. This regret bound was later improved for sparse systems in [IJR12]. [DMM+18, DMM+20] established regret and sample complexity results for the general stochastic LQR formulation with a polynomial dependence on natural system parameters. [FGKM18] studies the efficacy of first-order optimization-based approaches in synthesizing controllers for the LQR problem. The work of [MTR19] establishes that simple explore-then-commit strategies achieve the optimal dependence of regret on horizon, with [SF20] proving that such approaches are optimal in terms of their dependence on dimension of the system, as well. [CKM19] also gave a first optimal-in-horizon polynomial regret result simultaneously with the former work, albeit using an optimism-based approach. The result of [PH20] gives a rate-optimal regret bound for unknown systems with convex costs.
In terms of nonstochastic control, the method-of-moments system identification procedure and the corresponding analysis presented here appeared in [HKS20], which also gave the first end-to-end regret result for nonstochastic control starting with unknown system matrices. Even earlier, [SBR19] presented a regression procedure for identifying Markov operators under nonstochastic noise. Both of these approaches have roots in the classical Ho-Kalman system identification procedure [HK66], for which the first non-asymptotic sample complexity, although for stochastic noise, was furnished in [OO19]. For strongly convex costs, logarithmic regret was established in [Sim20]. Nonstochastic control of unknown, unstable systems without any access to prior information, also called black-box control, was addressed in [CH21].
@@ 8.7 Exercises
In previous chapters, we focused on control in dynamical systems, often assuming that the system state was fully observable and that system dynamics were either known or could be estimated through system identification. However, in many real-world scenarios, these assumptions do not hold. Instead, an agent may only have access to partial observations of the system and must learn to make accurate predictions without directly influencing the system’s evolution.
This chapter focuses on learning in a dynamical system, which means predicting future system outputs based on past observations while the underlying system parameters remain unknown. Unlike control, where an agent actively influences system trajectories to minimize a given cost, learning in this setting is a passive task; the goal is to accurately forecast future outputs without modifying the system’s behavior.
The ability to predict future states and outputs of an unknown system is fundamental to many domains, including time series forecasting, financial modeling, sensor networks, autonomous systems, and most recently large language models. Moreover, learning a predictive model is often a key stepping stone toward solving more complex problems, such as control under partial observability or reinforcement learning in dynamical systems. In this chapter, we study online learning algorithms for prediction in unknown linear dynamical systems, with an emphasis on regret minimization.
@@ 9.1 Learning in Dynamical Systems
The term learning in dynamical systems can refer to a variety of tasks, including state estimation, system identification, and policy learning. In this chapter, we focus on the specific task of prediction, where the objective is to estimate future outputs of a dynamical system based on past observations. Unlike classical system identification, which seeks to recover the system parameters explicitly, the learning approach taken here aims to optimize the predictions directly.
@@
For the remainder of this chapter, we consider a partially observable linear time-invariant dynamical system described by the following equations. The extension of the material in this chapter to time-varying linear dynamical systems is left as exercise 9.7.
@@
Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ is the observed control input, and $\mathbf{y}_{t}$ is the measured system output. The terms $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ account for the process noise and the observation noise, respectively. A crucial challenge in this setting is that the system matrices $A,B,C$ are unknown to the learner.
The learner’s objective is to construct an efficient prediction rule that estimates $\mathbf{y}_{t}$ based on past observations, despite not knowing the underlying parameters of the system.
@@ 9.2 Online Learning of Dynamical Systems
In an online learning setting, the learner sequentially observes the data and makes predictions in real time. Formally, the process unfolds as follows:
@@
    The learner makes a prediction $\hat{\mathbf{y}}_{t}$ for the next output.
@@
This regret-based framework allows us to compare different classes of predictors and design learning algorithms that efficiently adapt to unknown dynamical systems.
@@ 9.3 Classes of Prediction Rules
We now explore what constitutes a reasonable prediction rule for the output of a linear dynamical system. However, before considering the prediction rules themselves, we discuss how to compare classes of such rules. This is analogous to the definitions and discussion in Chapter 6 where we discuss comparison between policy classes for control. However, the definitions and the corresponding conclusions for making predictions are simpler.
@@
The proof of this lemma is left as an exercise to the reader at the end of this chapter. We now proceed to consider specific prediction rules.